OpenStudy (anonymous):

4[x-(3-2x)]+5=3(x+11)

OpenStudy (anonymous):

help

OpenStudy (whpalmer4):

\[4[x-(3-2x)]+5=3(x+11)\]\[4x-4(3-2x)+5=3x+33\]\[4x-12-(-8x)+5=3x+33\] \[4x+8x-12+5=3x+33\]Now subtract 3x from both sides and add 7 to both sides \[12x-3x-7+7=3x-3x+33+7\]\[9x=40\] \[x=40/9\]